OpenStudy (anonymous):

Find a solution to the first-order ODE: y'= e^(x-y), y(0)=1

myininaya (myininaya):

\[y'=e^xe^{-y}\] \[\frac{dy}{dx}=e^x e^{-y}\] \[e^{y} dy=e^{x} dx\]

myininaya (myininaya):

Can you do the rest ?

OpenStudy (anonymous):

Yes, thank you!! As soon as u typed that, i found it in my notes. There is another one I may need help with. Be on the look out pleeeease!!!

myininaya (myininaya):

k

OpenStudy (anonymous):

Im stuck at \[\int\ e ^{y} dy = \int\ e ^{x} dx\] I know that \[\int\limits_{?}^{?} e^x =e^x + C\] Can I do the same for e^y

OpenStudy (anonymous):

?

myininaya (myininaya):

\[\int\limits_{}^{} e^* d * =e^*+C\]

myininaya (myininaya):

It doesn't matter if * is x or y

myininaya (myininaya):

\[\int\limits_{}^{}e^y dy=e^y+k\]

OpenStudy (anonymous):

Oh, duh, lol. I keep making things more difficult. So ln(e^y+ k)= ln(e*+c) y= ln(e*+c)-k ?

myininaya (myininaya):

\[e^y+k=e^x+C\] \[e^y=e^x+C-k\] C-k is still a constant \[e^y=e^x+c\] I would leave it like this But if you are asked to solve for y Yes take ln( ) of both sides \[\ln(e^y)=\ln(e^x+c)\] \[y \ln(e)=\ln(e^x+c)\] \[y=\ln(e^x+c)\]

OpenStudy (anonymous):

Yes, im asked to solve for y, with y(0)=1. So 1= ln(e^(0)+c) I know that e^(0)=1, but ln(1)= 0.

myininaya (myininaya):

\[e^1=e^0+c => e=1+c => e-1=c\]

OpenStudy (anonymous):

Thaaaank yoou, oh kind one!!